solve by the substitution method 4x+6y=-4, -6x+y=26
The solution is??
(type an ordered pair)

Answer by ewatrrr(24785)


Hi
solve by the substitution method
-6x+y=26 Or y = 6x+26
4x+6y=-4
4x + 6(6x+26) = -4
40x = -160
x = -4 and y = 2
CHECKING our Answer***
-6*-4 + 2 = 24 + 2 = 26